Graph the demand curve for chiropractic visits using the points given in the table below. Answer...

  1. Graph the demand curve for chiropractic visits using the points given in the table below. Answer the questions below and interpret results. (What do results mean? Are they elastic? Inelastic?

P      Q

20    4

40    3

60    2

80     1

  • If price rises from $20 to 40, what is the price elasticity of demand at that point and what happens to the quantity demanded of chiropractic visits? Is this elastic or inelastic?
  • If price rises from $60 to $80?
  • $40 to 60?
Answer #1

By plotting the points we get the demand curve as shown below:

(i) If price rises from $20 to $40 ,quantity demanded for chiropractic visits decreases from 4 to 3 visits.

Price elasticity of demand = (% change in quantity demanded/ % change in price)

= [(3-4/4)] / [(40-20/ 20) ]

= (-1/4)/ (20/20)

= -1/4 = - 0.25.

Price elasticity of demand is inelastic because Ed<1 in absolute terms.

(ii) If price rises from $60 to $80 , then quantity demanded for chiropractic visits decreases from 2 to 1 visit.

Price elasticity of demand = (% change in quantity demanded/ % change in price)

= [(1-2/2)] / [(80-60/60)]

= (-1/2 )/ (20/60)

= -1/6 = - 0.167

Price elasticity of demand is inelastic because Ed<1 in absolute terms.

(iii) If price rises from $40 to $60 , then quantity demanded for chiropractic visits decreases from 3 to 2 visits.

Price elasticity of demand= (% change in quantity demanded/ % change in price)

= [(2-3/3)]/ [(60-40/40)]

= (-1/3)/ (20/40)

= -1/6 = -0.167

Price elasticity of demand is inelastic.